OpenStudy (anonymous):

Find the next three numbers in the pattern: 15, 21, 27, 33, 39 A. 45, 52, 59 B. 45, 51, 57 C. 44, 50, 56 D. 46, 52, 58

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (baru):

whats 15+6 ?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

They're multiplying by 3 but only the odds 3x5=15 3x7=21 3×9=27 3×11=33 3×13=39 3×15= 45 and so on

OpenStudy (michele_laino):

hint: you have an arithmetic sequence whose constant is \(d=6\) for example: \(21=15+6\) \(27=21+6\) \(33=27+6\) \(39=33+6\) \(39+6=...?\)

OpenStudy (anonymous):

B. 45, 51, 57. You add 6 each time.